OpenStudy (anonymous):

I need help with tis integral ln x/sqrt x dx we are using the method let u= and dv=

OpenStudy (amistre64):

so you are using integration by parts

OpenStudy (amistre64):

lnx is usually a sure bet for u

OpenStudy (amistre64):

1/sqrt(x) is simple enough to integrate into v

OpenStudy (amistre64):

you end up with a -int 2sqrt(x)/x which is powerable

OpenStudy (anonymous):

Yes we are using integration by parts. I found lots of examples but none that involved fractions so that is why I am having trouble.

OpenStudy (amistre64):

integrate x^(-1/2) what do you get?

OpenStudy (amistre64):

you should be comfortable with the power rule for integration

OpenStudy (anonymous):

-(1/2)-3/2

OpenStudy (amistre64):

\[\int x^n\to \frac{x^{n+1}}{n+1}\]

OpenStudy (anonymous):

x(-1/2+1)/ (-1/2+1)= x^0.5/0.5

OpenStudy (amistre64):

\[\binom{u=ln(x)}{du=\frac{1}{x}dx}\ \binom{v=2x^{1/2}}{dv=x^{-1/2}dx}\] \[\int udv=uv-\int vdu\]

OpenStudy (amistre64):

correct

OpenStudy (anonymous):

yes

OpenStudy (amistre64):

\[-\int vdu\to -\int 2x^{1/2}x^{-1}dx = -\int 2x^{-1/2} dx\]